OpenStudy (anonymous):

The probability of obtaining the sequence {H, T, T, H, H} when you toss a fair coin 5 times equals A. 1/32 B. 1/5 C. 1/2 D. 1/10 I think C but then again I suck at this..

OpenStudy (anonymous):

they want exactly that sequence {H, T, T, H, H} so {T, H, T, H, H} is not valid

OpenStudy (anonymous):

This is the problem I'm having it doesn't say whether the order is important in this case or not

OpenStudy (anonymous):

neigher is {H, H, T, T, H}

OpenStudy (anonymous):

surely they want us to recognize the notation.

OpenStudy (anonymous):

So you think it has to be in that exact order?

OpenStudy (anonymous):

do you know what it means when they use curved brackets?

OpenStudy (anonymous):

>So you think it has to be in that exact order? it sounds like it, I do not know the notation

OpenStudy (anonymous):

it is \((\frac{1}{2})^5\)

OpenStudy (anonymous):

but a sequence.......well what else than an exact sequence should it be?

OpenStudy (anonymous):

or if you prefer \[\frac{1}{2^5}\]

OpenStudy (anonymous):

you flip a coin 5 times, there are \(2^5=32\) equally likely elements in the sample space any particular sequence of head and tails has probability \(\frac{1}{32}\)

OpenStudy (anonymous):

@torontoXO the magic here is that the idiots give the exact sequence so it doesn't even matter how often you get a head or tail. they are only interested in their exact sequence. and of all the coins you can throw, there's just that one that gives everything like they have specified. that's one out of all possibilities.

OpenStudy (anonymous):

just one trail of events*

OpenStudy (anonymous):

basically if they said you need T-T-H then out of T-T-T T-T-H T-H-T T-H-H H-T-T H-T-H H-H-T H-H-H there's only one that is like they want. in your case you throw more often, so more possibilities how it can go.

OpenStudy (anonymous):

Ok so I think I get it. If it has to be in the exact order it only matters how often you could get the sequence or something like that?

OpenStudy (anonymous):

yes exactly you get the sequence, statistically, once with all the other possibilities if you want the highly specified then statistically you also get one of all the other with it.

OpenStudy (anonymous):

so if you throw coins like that 32 times, you have a chance of 1.0 to have the exact sequence in it. and if you only throw one time, well it's just 1/32 chance to get that one out of the 32.

OpenStudy (anonymous):

Ahhhhhhhh ok thanks a lot man hella appreciation